OpenStudy (anonymous):

Variable a is 4 more than variable b. Variable a is also 1 less than b. Which of the following pairs of equations best models the relationship between a and b? a = b + 4 a = b − 1 a = 4b a = b + 1 b = a + 4 b = a − 1 b = 4a b = a + 1

OpenStudy (johnweldon1993):

So we have 'a' and 'b' If 'a' is 4 more than 'b' ...that is just addition right? a = b + 4 And variable 'a' is 1 less than 'b' so a = b - 1

OpenStudy (anonymous):

wait im confused lol, cause its two equations for each answer so im confused rn

OpenStudy (johnweldon1993):

Right which is why we needed to work with them both if A is 4 more than B...then B + 4 = A *****************************************First equation we needed and if A is 1 less than B...then B-1 = A *****************************************Second equation we needed So our answer choice with \(\large a = b + 4\) and \(\large a = b - 1\) would be correct

OpenStudy (anonymous):

thank you so muchh

OpenStudy (johnweldon1993):

Not a problem :) Enjoy your summer!